# MA-RLHF: Reinforcement Learning from Human Feedback with Macro Actions

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) has demonstrated effectiveness in aligning large language models (LLMs) with human preferences. However, token-level RLHF suffers from the credit assignment problem over long sequences (Bengio et al., 2013), where delayed rewards make it challenging for the model to discern which actions contributed to successful outcomes. This hinders learning efficiency and slows convergence (Mann & Mannor, 2014; Machado et al., 2023). In this paper, we propose MA-RLHF, a simple yet effective RLHF framework that incorporates macro actions— sequences of tokens or higher-level language constructs—into the learning process. By operating at this higher level of abstraction, our approach reduces the temporal distance between actions and rewards, facilitating faster and more accurate credit assignment. This results in more stable policy gradient estimates and enhances learning efficiency within each episode, all without increasing computational complexity during training or inference. We validate our approach through extensive experiments across various model sizes and tasks, including text summarization, dialogue generation, question answering, and program synthesis. Our method achieves substantial performance improvements over standard RLHF, with performance gains of up to 30% in text summarization and code generation, 18% in dialogue, and 8% in question answering tasks. Notably, our approach reaches parity with vanilla RLHF 1.7x to 2x faster in terms of training time and continues to outperform it with further training. We will make our code and data publicly available at https://github.com/ernie-research/MA-RLHF.
@@ 1 Introduction
Existing RLHF (Ouyang et al., 2022; Bai et al., 2022; Askell et al., 2021) methods mainly optimize decisions at the level of individual tokens, and require to process a vast number of minute adjustments. However, this fine-grained training paradigm can lead to the credit assignment problem (Bengio et al., 2013; Kaelbling et al., 1996), particularly when dealing with long-distance dependencies. As LLM agents attempt to optimize decisions across extensive sequences, the difficulty in attributing the credits of actions to specific tokens complicates the reinforcement learning (RL) process (Pignatelli et al., 2024). Moreover, the use of subword tokenization, such as Byte-Pair Encoding (Sennrich et al., 2016), often splits words into smaller pieces. For instance, OpenAI’s ChatGPT [^1] https://platform.openai.com/tokenizer treats each token as three quarters of a word on average, resulting in sequences that are 33% longer than word counts (OpenAI, 2024) and further exacerbates the credit assignment problem.
Additionally, standard RLHF methods may overlook essential linguistic structures and dependencies between adjacent tokens in natural language. For example, consider the phrase Big Apple [^2] https://en.wikipedia.org/wiki/Big_Apple, treating Big and Apple as isolated decisions misses the cohesive meaning of the term, which actually refers to the “New York City”. The token-level granularity of natural language can hinder the agent’s ability to capture high-level language constructs in RL optimization, as some sequences are better understood when evaluated holistically.
To address these challenges, we propose a new framework called macro-action RLHF (MA-RLHF) that incorporate macro action — sequences of tokens or high-level language constructs — into the RLHF framework. The concept of macro actions, has been explored in the literature of planning  (Iba, 1989; Korf, 1985; Sacerdoti, 1974) and reinforcement learning (Thrun & Schwartz, 1994; Precup et al., 1997; Hauskrecht et al., 2013), simplifies decision-making by operating at high levels of temporal abstraction under the framework of semi-Markov Decision Processes (SMDPs) (Sutton et al., 1999). Macro actions leverage meaningful linguistic structures and allow the agent to learn from “long-sighted” macro-level actions instead of “short-sighted” token-level actions, which can potentially leads to improved learning efficiency and scalability. Alternatively, MA-RLHF can also be interpreted from the perspective of reversing tokenization; MA-RLHF serves as a de-tokenization process to reconstruct high-level language units from subword pieces. By merging tokens into macro actions, we reduce the number of decision points and shorten decision trajectories, alleviating the credit assignment problem caused by long temporal distances.
@@
  • We propose MA-RLHF, a simple yet effective RLHF framework that integrate the macro actions into RLHF to align LLMs with human preference. We demonstrate the effectiveness of our approach through extensive experiments across various datasets and tasks, including text summarization, dialogue generation, question answering, and program synthesis.
  • We show that MA-RLHF achieves 1.7x to 2x faster convergence in reward scores during training compared to the standard token-level RLHF, without introducing additional computational costs during training or inference. MA-RLHF also exhibits strong scalability across model sizes ranging from 2B to 27B parameters.
@@ 2.1 Reinforcement Learning and Policy Optimization
Problem Definition   RL addresses the problem of finding an policy to make optimal sequential decisions in environments modeled as the Markov Decision Processes (MDPs) (Sutton & Barto, 1999). An MDP is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\rho_{0},\gamma)$, where $\mathcal{S}$ denotes a finite set of states, $\mathcal{A}$ is a finite set of actions, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the state transition probability distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\rho_{0}:\mathcal{S}\rightarrow[0,1]$ defines the initial state distribution, and $\gamma\in(0,1)$ is the discount factor that determines the importance of future rewards.
@@
The state-action value function $Q_{\pi}(s_{t},a_{t})=\mathbb{E}_{s_{t+1},a_{t+1},\dots}\left[\sum_{l=0}^{\infty}\gamma^{l}r(s_{t+l})\right]$ measures the expected return of taking action $a_{t}$ at state $s_{t}$ and following policy $\pi$ thereafter. The value function $V_{\pi}(s_{t})=\mathbb{E}_{a_{t},s_{t+1},\dots}\left[\sum_{l=0}^{\infty}\gamma^{l}r(s_{t+l})\right]$ estimates the expected return from state $s_{t}$ under the policy $\pi$. The advantage function $A_{\pi}(s_{t},a_{t})=Q_{\pi}(s_{t},a_{t})-V_{\pi}(s_{t})$ reflects the relative value of taking action $a_{t}$ at state $s_{t}$ compared to the average value of the state.
@@
The goal of RL is to find an optimal policy $\pi_{\theta}(a\mid s)$, parameterized by $\theta$, that maximizes the expected cumulative discounted reward: $J(\theta)=\mathbb{E}_{s_{0},a_{0},\dots}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t})\right]$, where $s_{0}\sim\rho_{0}(s_{0})$ represents the initial state distribution, $a_{t}\sim\pi_{\theta}(a_{t}\mid s_{t})$ denotes the action selection based on the policy, and $s_{t+1}\sim P(s_{t+1}\mid s_{t},a_{t})$ specifies the state transition dynamics.
@@
Proximal Policy Optimization   Policy gradient methods are a common approach for optimizing policies by estimating the gradient of a performance objective with respect to the policy parameters $\theta$. The policy gradient is given by: $\nabla_{\theta}J(\theta)=\mathbb{E}\left[\sum_{t=0}^{\infty}A_{t}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\right]$, where the expectation $\mathbb{E}[\cdot]$ is taken over the randomness of initial state, policy, and state-transition. The policy gradient guides us how to adjust the policy parameters to improve the expected return. Among the family of policy gradient methods, Proximal Policy Optimization (Schulman et al., 2017, PPO) is perhaps the most widely-used one due to its simplicity and empirical effectiveness. PPO simplifies TRPO (Schulman et al., 2015) by using a clipped surrogate objective function to penalize large deviations from the old policy, thereby ensuring more stable updates. Specifically, PPO introduces a clipped objective function:
@@
$J^{\text{ppo-clip}}(\theta)=\mathbb{E}\left[\min\left(\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}{A}_{t},\text{clip}(\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})},1-\epsilon,1+\epsilon){A}_{t}\right)\right],$ (1)
@@
where $\epsilon$ is a hyperparameter that defines the range for clipping. Nowadays, PPO usually comes as the first choice for RL practitioners.
@@ 2.2 RLHF for Human Alignment
(1) SFT stage: A pre-trained language model (LM) is fine-tuned on a dataset of human demonstrations, learning to generate responses that align with human instructions and preferences.
(2) RM stage: A reward model $r_{\phi}(x,y)$, parameterized by $\phi$, is trained to predict human preference scores for model outputs. The training data $\mathcal{D}={(x_{i},y_{i}^{+},y_{i}^{-})}_{i=1}^{N}$ consists of prompts $x_{i}$ and pairs of responses $(y_{i}^{+},y_{i}^{-})$, where $y_{i}^{+}$ is preferred over $y_{i}^{-}$ by human annotators. The reward model is trained using the ranking loss: $\mathcal{L}_{\textrm{RM}}=-\log\sigma(\log(r_{\phi}(x,y_{+})-r_{\phi}(x,y_{-})))$, where $\sigma$ denotes the sigmoid function, and $r_{\phi}$ represents the reward function parameterized by $\phi$.
@@
$R(x,y)=r_{\phi}(x,y)+\beta D_{\mathrm{KL}}(\pi_{\mathrm{sft}}(y~{}|~{}x)~{}||~{}\pi_{\theta}(y~{}|~{}x)),$
@@
In the RLHF stage, the PPO algorithm, as detailed in eq. 1, is employed to optimize the RL policy. In the context of RLHF, we denote the state $s_{t}=\{a_{1},a_{2},\dots,a_{t-1}\}$ as the sequence of tokens generated up to time step $t$, while $a_{t}$ represents the token selected at the $t$-th position.
@@ 3.1 Revisiting Macro Actions (Options)
Macro actions, also referred to as options (Sutton et al., 1999), are high-level constructs that encapsulate a sequence of primitive actions (i.e., subword tokens); by its definition, macro actions allows an agent to operate at a coarser temporal scale.
Formally, a macro action is characterized by three components: (1) a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ which guides the action selection among macro actions; (2) a termination condition $\zeta:\mathcal{S}^{+}\rightarrow[0,1]$, which determines where the macro action should end; (3) a initiation set $\mathcal{I}\subseteq\mathcal{S}$, which is a subset of states that macro actions can begin with. Once initiated with a state $s_{0}\in\mathcal{I}$, the macro action follows policy $\pi$ until it reaches the termination condition according to $\zeta$. Intuitively, the use of carefully designed macro actions can extend decision-making temporally, it allows the agent to avoid “short-sighted” token-level decisions and encourage “long-sighted” macro-level decisions, thereby simplifies the decision-making process and potentially enhances learning efficiency.
@@ 3.2 RLHF with Macro Actions
Figure 1: Illustration of the MA-RLHF framework. Standard RLHF makes decisions and evaluates value scores at the token level, while MA-RLHF operates at the macro action level, making decisions over sequences of tokens at a coarser temporal scale.
@@ 3.2.1 Formalization of Macro Actions
  • Perplexity-based (PPL) termination: Perplexity measures the likelihood of a sequence of tokens. Here, the perplexity of a macro action is proportional to the averaged entropy of the token within it, $i.e.$, $\mathrm{ppl}(\omega_{\tau})\propto-\frac{1}{|\omega_{\tau}|}\sum_{a\in\omega_{\tau}}\log p_{a}$. A macro action terminates until it reaches a token that has negative impact on the perplexity of the macro action. Mathematically, we construct $\omega_{\tau}=\{a_{t_{\tau}},\ldots,a_{t_{\tau+1}-1}\}$ such that $\mathrm{ppl}(\omega_{\tau}\cup a_{\tau+1})>\mathrm{ppl}(\omega_{\tau})$ and $\mathrm{ppl}(\{a_{\tau},\ldots,a_{i}\})\geq\mathrm{ppl}(\{a_{\tau},\ldots,a_{i+1}\})$ for all $t_{\tau}\leq i\leq t_{\tau+1}$.
@@ 3.2.2 Policy Optimization with Macro Actions
In MA-RLHF, we adapt the PPO algorithm to operate at the macro action level, referred to as MA-PPO. The policy $\pi_{\theta}$ still outputs probabilities over individual tokens, but for optimization, we consider the joint probability of the macro action: $\pi_{\theta}(\omega_{\tau}~{}|~{}s_{\tau})=\prod_{t=t_{\tau}}^{t_{\tau+1}}\pi_{\theta}(a_{t}~{}|~{}a_{<t})$. The macro reward for executing the macro action $\omega_{\tau}$ at the macro time step $\tau$ is defined as: $R_{\tau}=\mathbb{E}\big{[}\sum_{i=0}^{|\omega_{\tau}|-1}\rho^{i}r_{t_{\tau}+i}~{}\big{|}~{}s_{\tau}\big{]},$ where $r_{t}$ is the reward received at time step $t$, and we set the discount factor $\rho=1$ in our experiments.
@@
$\mathcal{L}^{\text{MA-PPO}}(\theta)={\mathbb{E}}_{\tau}\left[\min\left(r_{\tau}(\theta)\hat{A}_{\tau},\ \text{clip}\left(r_{\tau}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{\tau}\right)\right],$
where $r_{\tau}(\theta)=\pi_{\theta}\big{(}\omega_{\tau}\mid s_{\tau}\big{)}/\pi_{\theta_{\text{old}}}\big{(}\omega_{\tau}\mid s_{\tau}\big{)}$, and $\hat{A}_{\tau}$ is the estimated advantage at macro time step $\tau$, $\epsilon$ is a constant that defines the range for clipping, and $\pi_{\theta_{\text{old}}}$ is the policy before the update.

### v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) has demonstrated effectiveness in aligning large language models (LLMs) with human preferences. However, token-level RLHF suffers from the credit assignment problem over long sequences, where delayed rewards make it challenging for the model to discern which actions contributed to preferred outcomes. This hinders learning efficiency and slows convergence. In this paper, we propose MA-RLHF, a simple yet effective RLHF framework that incorporates macro actions — sequences of tokens or higher-level language constructs — into the learning process. By operating at higher level of abstraction, our approach reduces the temporal distance between actions and rewards, facilitating faster and more accurate credit assignment. This results in more stable policy gradient estimates and enhances learning efficiency within each episode, all without increasing computational complexity during training or inference. We validate our approach through extensive experiments across various model sizes and tasks, including text summarization, dialogue generation, question answering, and program synthesis. Our method achieves substantial performance improvements over standard RLHF, with performance gains of up to 30% in text summarization and code generation, 18% in dialogue, and 8% in question answering tasks. Notably, our approach reaches parity with vanilla RLHF $1.7\sim 2$ times faster in terms of training time and continues to outperform it with further training. We make our code and data publicly available at https://github.com/ernie-research/MA-RLHF.
@@ 1 Introduction
Existing RLHF (Ouyang et al., 2022; Bai et al., 2022; Askell et al., 2021) methods mainly optimize decisions at the level of individual tokens, and require to process a vast number of minute adjustments. However, this fine-grained training paradigm can lead to the credit assignment problem (Kaelbling et al., 1996; Pang et al., 2019; Machado et al., 2023b; Pignatelli et al., 2023), particularly when dealing with long-distance dependencies. As LLM agents attempt to optimize decisions across extensive sequences, the difficulty in attributing the credits of actions to specific tokens complicates the reinforcement learning (RL) process (Pignatelli et al., 2024). Moreover, the use of subword tokenization, such as Byte-Pair Encoding (Sennrich et al., 2016), often splits words into smaller pieces. For instance, OpenAI’s ChatGPT [^1] https://platform.openai.com/tokenizer treats each token as three quarters of a word on average, resulting in sequences that are 33% longer than word counts (OpenAI, 2024) and further exacerbates the credit assignment problem.
Additionally, standard RLHF methods may overlook essential local co-occurrence patterns or inherent structures between adjacent tokens in natural language. For example, consider the phrase Big Apple [^2] https://en.wikipedia.org/wiki/Big_Apple, treating Big and Apple as isolated decisions misses the cohesive meaning of the term, which actually refers to the “New York City”. The token-level granularity of natural language can hinder the agent’s ability to capture high-level language constructs in RL optimization, as some sequences are better understood when evaluated holistically.
To address these challenges, we propose a new framework called macro-action RLHF (MA-RLHF) that incorporate macro action — sequences of tokens or high-level language constructs — into the RLHF framework. The concept of macro actions, has been explored in the literature of planning  (Iba, 1989; Korf, 1985; Sacerdoti, 1974) and reinforcement learning (Thrun & Schwartz, 1994; Precup et al., 1997; Hauskrecht et al., 2013), simplifies decision-making by operating at high levels of temporal abstraction under the framework of semi-Markov Decision Processes (SMDPs) (Sutton et al., 1999b). Macro actions leverage temporal abstraction by chunking the sequences and reducing the decision resolution, enabling the agent to learn from “long-sighted” macro-level actions instead of “short-sighted” token-level actions. This can potentially lead to improved learning efficiency and scalability. Alternatively, MA-RLHF can also be interpreted from the perspective of reversing tokenization; MA-RLHF serves as a de-tokenization process to reconstruct high-level language units from subword pieces. By merging tokens into macro actions, we reduce the number of decision points and shorten decision trajectories, alleviating the credit assignment problem caused by long temporal distances.
@@
  • We propose MA-RLHF, a simple yet effective RLHF framework that integrates the macro actions into RLHF to align LLMs with human preference. We demonstrate the effectiveness of our approach through extensive experiments across various datasets and tasks, including text summarization, dialogue generation, question answering, and code generation.
  • We show that MA-RLHF achieves 1.7$\times$ to 2$\times$ faster learning efficiency in reward scores during training compared to the standard token-level RLHF, without introducing additional computational costs during training or inference. MA-RLHF also exhibits strong scalability across model sizes ranging from 2B to 27B parameters.
@@ 2.1 Reinforcement Learning and Policy Optimization
Problem Definition   RL addresses the problem of finding a policy to make optimal sequential decisions in environments modeled as a Markov Decision Process (MDP) (Sutton & Barto, 1999). An MDP is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\rho_{0},\gamma)$, where $\mathcal{S}$ denotes a finite set of states, $\mathcal{A}$ is a finite set of actions, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the state transition probability distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\rho_{0}:\mathcal{S}\rightarrow[0,1]$ defines the initial state distribution, and $\gamma\in(0,1)$ is the discount factor that determines the importance of future rewards.
@@
Given a trajectory $(s_{0},a_{0},s_{1},a_{1},\cdots)$, a reward $r_{t}=r(s_{t},a_{t})$ is received at each time $t$. The state-action value function $Q_{\pi}(s_{t},a_{t})=\mathbb{E}_{s_{t+1},a_{t+1},\dots}\left[\sum_{l=0}^{\infty}\gamma^{l}r_{t+l}\right]$ measures the expected return of taking action $a_{t}$ at state $s_{t}$ and following policy $\pi$ thereafter. The value function $V_{\pi}(s_{t})=\mathbb{E}_{a_{t},s_{t+1},\dots}\left[\sum_{l=0}^{\infty}\gamma^{l}r_{t+l}\right]$ estimates the expected return from state $s_{t}$ under the policy $\pi$. The advantage function $A_{\pi}(s_{t},a_{t})=Q_{\pi}(s_{t},a_{t})-V_{\pi}(s_{t})$ reflects the relative value of taking action $a_{t}$ at state $s_{t}$ compared to the average value of the state.
@@
The goal of RL is to find an optimal policy $\pi_{\theta}(a\mid s)$, parameterized by $\theta$, that maximizes the expected cumulative discounted reward: $J(\theta)=\mathbb{E}_{s_{0},a_{0},\dots}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $s_{0}\sim\rho_{0}(s_{0})$ represents the initial state distribution, $a_{t}\sim\pi_{\theta}(a_{t}\mid s_{t})$ denotes the action selection based on the policy, and $s_{t+1}\sim P(s_{t+1}\mid s_{t},a_{t})$ specifies the state transition dynamics.
@@
Proximal Policy Optimization   Policy gradient methods are a common approach for optimizing policies by estimating the gradient of a performance objective with respect to the policy parameters $\theta$. The policy gradient is given by: $\nabla_{\theta}J(\theta)=\mathbb{E}\left[\sum_{t=0}^{\infty}A_{t}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\right]$, where the expectation $\mathbb{E}[\cdot]$ is taken over the randomness of the initial state, policy, and state-transition. The policy gradient guides us how to adjust the policy parameters to improve the expected return. Among the family of policy gradient methods, Proximal Policy Optimization (Schulman et al., 2017, PPO) is perhaps the most widely-used one due to its simplicity and empirical effectiveness. PPO simplifies TRPO (Schulman et al., 2015) by using a clipped surrogate objective function to penalize large deviations from the old policy, thereby ensuring more stable updates. Specifically, PPO introduces a clipped objective function:
@@
$J^{\text{ppo-clip}}(\theta)=\mathbb{E}_{t}\left[\min\left(\frac{\pi_{\theta}(a_{t}\!\mid\!s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\!\mid\!s_{t})}{A}_{t},\text{clip}(\frac{\pi_{\theta}(a_{t}\!\mid\!s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\!\mid\!s_{t})},1-\epsilon,1+\epsilon){A}_{t}\right)\right],$ (1)
@@
where $\epsilon$ is a hyperparameter that defines the range for clipping. The expectation $\mathbb{E}_{t}[\dots]$ indicates the empirical average over a finite batch of samples. Nowadays, PPO usually comes as the first choice for RL practitioners.
@@ 2.2 RLHF for Human Alignment
(1) Supervised Fine-Tuning (SFT) stage: A pre-trained language model (LM) is fine-tuned on a dataset of human demonstrations, learning to generate responses that align with human instructions and preferences.
(2) Reward Modeling (RM) stage: A reward model is trained on a labeled preference dataset $\mathcal{D}={(x_{i},y_{i}^{+},y_{i}^{-})}_{i=1}^{N}$, consisting of prompts $x_{i}$ and pairs of responses $(y_{i}^{+},y_{i}^{-})$, where $y_{i}^{+}$ is preferred over $y_{i}^{-}$ by human annotators. The reward model $r_{\phi}(x,y)$, parameterized by $\phi$, is trained using the ranking loss: $\mathcal{L}_{\textrm{RM}}=-\log\sigma(\log(r_{\phi}(x,y_{+})-r_{\phi}(x,y_{-})))$, where $\sigma$ denotes the sigmoid function.
@@
$R(x,y)=r_{\phi}(x,y)-\beta D_{\mathrm{KL}}(\pi_{\theta}(\cdot\mid x)~{}\|~{}\pi_{\mathrm{sft}}(\cdot\mid x)),$
@@
In the RLHF stage, the PPO algorithm, as detailed in Equation 1, is employed to optimize the RL policy. In the context of RLHF, we denote the state $s_{t}=\{s_{0},a_{0},a_{1},\dots,a_{t-1}\}$ as the sequence of tokens generated up to time step $t$, while $s_{0}$ represents the initial states, i.e., the prompt, and $a_{t}$ represents the token selected at the $t$-th position.
@@ 3.1 Revisiting Macro Actions (Options)
Macro actions, also referred to as options (Sutton et al., 1999b), are high-level constructs that encapsulate a sequence of primitive actions (i.e., subword tokens); by its definition, macro actions allows an agent to operate at a coarser temporal scale.
Formally, a macro action is characterized by three components: (1) a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ which guides the action selection among actions; (2) a termination condition $\zeta:\mathcal{S}^{+}\rightarrow[0,1]$, which determines where the macro action should end; (3) a initiation set $\mathcal{I}\subseteq\mathcal{S}$, which is a subset of states that macro actions can begin with. Once initiated with a state $s_{0}\in\mathcal{I}$, the macro action follows policy $\pi$ until it reaches the termination condition according to $\zeta$. Intuitively, the use of carefully designed macro actions can extend decision-making temporally, it allows the agent to avoid “short-sighted” token-level decisions and encourage “long-sighted” macro-level decisions, thereby simplifies the decision-making process and potentially enhances learning efficiency.
@@ 3.2 RLHF with Macro Actions
Figure 1: Illustration of the MA-RLHF optimization framework. Standard RLHF makes decisions and evaluates value scores at the token level, while MA-RLHF makes decisions over sequences of tokens at a coarser temporal scale.
@@ 3.2.1 Formalization of Macro Actions
  • Perplexity-based (PPL) termination: Perplexity measures the likelihood of a sequence of tokens. Here, the perplexity of a macro action is proportional to the averaged entropy of the token within it, $i.e.$, $\mathrm{ppl}(\omega_{\tau})\propto-\frac{1}{|\omega_{\tau}|}\sum_{a\in\omega_{\tau}}\log p_{a}$. A macro action terminates until it reaches a token that has negative impact on the perplexity of the macro action. Mathematically, we construct $\omega_{\tau}=\{a_{t_{\tau}},\ldots,a_{t_{\tau+1}-1}\}$ such that $\mathrm{ppl}(\omega_{\tau}\cup a_{t_{\tau+1}})>\mathrm{ppl}(\omega_{\tau})$ and $\mathrm{ppl}(\{a_{t_{\tau}},\ldots,a_{i}\})\geq\mathrm{ppl}(\{a_{t_{\tau}},\ldots,a_{i+1}\})$ for all $t_{\tau}\leq i\leq t_{\tau+1}-2$.
After determining the macro action based on the termination condition, we apply the state value function and importance sampling at the macro level Equation 1. We provide the details of implementation in Section D.1.
@@ 3.2.2 Policy Optimization with Macro Actions
In MA-RLHF, we adapt the PPO algorithm for optimization, referred to as MA-PPO. In the context of LLMs, expanding the action space with additional macro actions/tokens results in re-architecting the LLM’s vocabulary and retraining the model, which is computationally prohibitive. Thus, we maintain the original action space as pretrained LLMs, which can be treated as “single-step” primitive options as noted in (Sutton et al., 1999b). The policy $\pi_{\theta}$ still outputs probabilities over individual tokens, but for optimization, we consider the joint probability of the macro action: $\pi_{\theta}(\omega_{\tau}~{}|~{}s_{\tau})=\prod_{t=t_{\tau}}^{t_{\tau+1}}\pi_{\theta}(a_{t}~{}|~{}a_{<t})$. The macro reward for executing the macro action $\omega_{\tau}$ at the macro time step $\tau$ is defined as: $R_{\tau}=\mathbb{E}\big{[}\sum_{i=0}^{|\omega_{\tau}|-1}\rho^{i}r_{t_{\tau}+i}~{}\big{|}~{}s_{\tau}\big{]},$ where $r_{t}$ is the reward received at time step $t$, and we set the discount factor $\rho=1$ in our experiments.
@@
$\mathcal{L}^{\text{MA-PPO}}(\theta)={\mathbb{E}}_{\tau}\left[\min\left(\frac{\pi_{\theta}\big{(}\omega_{\tau}\mid s_{\tau}\big{)}}{\pi_{\theta_{\text{old}}}\big{(}\omega_{\tau}\mid s_{\tau}\big{)}}\hat{A}_{\tau},\ \text{clip}\left(\frac{\pi_{\theta}\big{(}\omega_{\tau}\mid s_{\tau}\big{)}}{\pi_{\theta_{\text{old}}}\big{(}\omega_{\tau}\mid s_{\tau}\big{)}},1-\epsilon,1+\epsilon\right)\hat{A}_{\tau}\right)\right],$
where $\hat{A}_{\tau}$ is the estimated advantage at macro time step $\tau$, $\epsilon$ is a constant that defines the range for clipping, and $\pi_{\theta_{\text{old}}}$ is the policy before the update.